# Reducing the Cost of Energy Differences in Variational Monte Carlo with Spotlight Sampling

### v1

@@ Abstract
We investigate an approximate sampling scheme that can significantly reduce the cost scaling of variational Monte Carlo when it is employed to predict the energy differences associated with local chemical changes. Inspired by side-chaining and embedding methods, this spotlight sampling approach adopts an approximate fragmented Hamiltonian and correlated sampling to reduce cost scaling to the point that it is essentially linear with system size, with the potential to go sub-linear if certain conditions are met. In tests on bond stretching energies in alcohols and hydrogen dimer chains, we observe the anticipated linear scaling as well as an explicit cost crossover with standard variational Monte Carlo.
preprint: AIP/123-QED
I Introduction
Simultaneously capturing the effects of both weak and strong electron correlation remains challenging for traditional quantum chemistry methods, especially in excited states and large systems. Benjamin G. Levine and Martínez (2006); Sherrill (2005); Hait et al. (2019); Aoto et al. (2017); Van Voorhis and Head-Gordon (2000); Harvey (2006); Vogiatzis et al. (2017); Olsen (2011) In many difficult cases, Quantum Monte Carlo (QMC) methods have proven to be effective alternatives. Shee et al. (2019); Needs et al. (2020); Kent et al. (2020); Clark et al. (2011); Assaraf, Moroni, and Filippi (2017); Petruzielo, Toulouse, and Umrigar (2012); Emmanuel Giner and Toulouse (2016); Dash et al. (2019, 2021); Pineda Flores and Neuscamman (2019); Otis and Neuscamman (2023); Otis, Craig, and Neuscamman (2020) For example, in some excited states, even variational Monte Carlo (VMC), perhaps the most straightforward QMC technique, has proven more accurate than full-triples equation of motion coupled cluster theory. Otis and Neuscamman (2023); Otis, Craig, and Neuscamman (2020) However, the significant computational cost required to evaluate large numbers of random samples continues to limit the applicability of both VMC and QMC methods more generally. Indeed, such sampling places an unusually large prefactor in front of VMC’s quartic cost growth with system size, Bienvenu et al. (2022) meaning that the advantages of this modest polynomial scaling can be difficult to realize in practice. In this study, we show that a combination of approximate fragmentation and correlated sampling can reduce both cost and scaling while maintaining accuracy when using VMC to predict energy differences associated with local chemical changes, as for example when stretching a chemical bond.
It has long been recognized that, thanks to correlated sampling techniques, it can be advantageous to directly estimate energy differences in VMC rather than attempting to make independent estimates of the energy before and after a chemical change. Umrigar (1989) The basic idea is that, for regions of the system that are only weakly affected by the change, statistical fluctuations will mostly cancel out so long as essentially the sampling of electron positions in those regions is used on either side of the energy difference. The space warp transformation, which we employ in the present study, achieves this goal explicitly. Sorella and Capriotti (2010); Filippi and Umrigar (2000) What’s more, by trading the task of estimating two extensive energies for the task of estimating an intensive energy difference, such correlated sampling approaches should in general drop the cost scaling of VMC from quartic to cubic. Here, we seek to combine this advantage with the advantages of local fragmentation approaches, in particular approaches inspired by recent side-chaining methods. Bienvenu et al. (2022); Feldt and Assaraf (2021)
Locality has, more generally, long been exploited in attempts to mitigate the cost of VMC. Kohn (1959); Maslen et al. (1998); Burant, Scuseria, and Frisch (1996); Schwegler and Challacombe (1996) Early works focused on leveraging the sparsity of the Slater matrix when using local orbitals in order to reduce the costs associated with Slater determinant ratios, Williamson, Hood, and Grossman (2001); Manten and Lüchow (2003); Alfè and Gillan (2004); Austin et al. (2006) which can also be mitigated by complimentary delayed update methods. McDaniel et al. (2017) More recently, Feldt and coworkers have shown that that evaluating local side chains, in which only a few of the electrons near a given nucleus are allowed to move, is an effective way to reduce the cost of dealing with core electrons. Feldt and Assaraf (2021) This side-chaining approach was then generalized in a study by Bienvenu and coworkers, in which they show that it can reduce the cost of VMC sampling by one power of the system size. Bienvenu et al. (2022)
These side-chaining methods have strong parallels to embedding methods, Knizia and Chan (2012); Sun and Chan (2016); Welborn, Tsuchimochi, and Van Voorhis (2016); Ye et al. (2019) in that the side chains sample some electrons well while not sampling the “environment” electrons at all, but they differ in one crucial respect. Unlike embedding methods, in which the interaction energy between the system and its environment is approximated, these two side-chain studies carefully set up their overall energy estimates to be unbiased, and so they do not add any approximation. This juxtaposition begs the question: what might we be able to gain if we instead intentionally introduced simplifications that biased the results? Can the cost growth with system size be reduced further while maintaining reasonable accuracy for relevant chemical quantities like bond stretching energies? In this study, we embark on a preliminary investigation of these questions, concluding that, yes, for some types of predictions, a combination of correlated sampling, side-chaining-inspired fragmentation, and embedding-inspired Hamiltonian approximations allow for VMC cost scaling to be reduced further, in principle to the point that it can become linear or even sub-linear in the system size.
This paper is organized as follows. We begin in Section II.1 with a theoretical analysis of how the cost scaling of drawing samples in VMC can be reduced by a combination of various techniques, within which we include a review of the fast Slater determinant math used in Markov chains in which only a small number of electrons move. During this analysis, we do not address the question of accuracy, relying instead on the assumption that certain sampling conditions can be met without doing too much damage to VMC’s energy difference predictions. After completing that analysis, we then introduce in Section II.2 an explicit approximation scheme, spotlight sampling, which seeks to meet these conditions while maintaining accuracy, focusing in particular on how to deal with both erroneous Pauli exclusion effects, and, in Section II.3, medium and long range Coulomb interactions. With an approximate Hamiltonian defined, Section II.4 then completes the cost analysis by showing that its local energy evaluations do not alter the conclusions of the initial sampling cost analysis. We then touch on (Section II.5) how this approach intersects with the more general VMC challenge of maintaining numerical stability during long chains of low-rank Slater matrix inverse updates before turning to our results. In these results (Section III), we first verify the necessity of spotlight sampling’s buffering approach, after which we explore proof of principle calculations that demonstrate the cost advantages and test accuracy in a series of alcohols’ O-H bond stretches. We then end in Section IV with some concluding remarks.
II Theory
II.1 Reducing the Cost of Sampling
@@
To reduce the cost of working with $M^{-1}$, it is updated at $O(N^{2})$ cost via the Sherman-Morrison formula after each accepted move. Ceperley (1978); Umrigar (1993); Lee et al. (2011) Generating a new and independent sample for use in Eq. (1) requires $O(N)$ of these one-electron moves, and, at least when we are attempting to estimate the total energy $E$, the overall number of samples $K$ needs to grow linearly with $N$ to counteract the linear growth in the variance. Bienvenu et al. (2022) All told, the computational effort comes out as $O(N^{4})$, which breaks down as taking $O(N)$ samples, each of which requires $O(N)$ moves, each of which bears an $O(N^{2})$ cost for the Sherman-Morrison update.
In this study, we will explore a combination of correlated sampling and fragmentation in order to approximate energy differences for local chemical changes at reduced cost. To start, let us focus first on the effort required to draw samples, and in particular how a fragmentation of the system can reduce this effort. Imagine if, as in a PMC side chain, Bienvenu et al. (2022) we only ever had to consider situations in which a small group of O(1) electrons had moved away from their initial positions, while all other electrons remained frozen where they started. In that case, only O(1) of the rows of $M$ would ever change, allowing for an efficient evaluation of a key determinant ratio.
@@
In this study, our goal is more specific than total energies: by also leveraging correlated sampling, we wish to predict energy changes due to local chemical changes at an even lower cost. To do so, we make use of Filippi and Umrigar’s space warp correlated sampling technique. Filippi and Umrigar (2000) In this approach, whose mathematical details are in the Appendix, electron positions are sampled at one molecular geometry and “warped” to the other geometry in a way that leaves unchanged the positions of electrons near atoms that did not move. We would therefore anticipate that the energy difference $\Delta E^{(l)}$ and energy difference uncertainty $\sigma^{(l)}$ for fragment $l$ will be relatively small if the $l$th fragment’s atoms do not move during the local change, and relatively large if its atoms do move. If we were to sort the fragments in decreasing order of $\sigma^{(l)}$ and could bound these uncertainties by $a/l$ when using some fixed and $O(1)$ fragment sample size $K_{f}$, where $a$ is some constant, then estimating the overall energy difference for the local change would no longer require us to take $O(N)$ samples per fragment. Indeed, $O(1)$ samples per fragment would now suffice. For example, with our bound, $K_{f}$ samples per fragment would give an overall uncertainty of
@@ II.2 Spotlight Sampling
II.3 Long Range Electrostatics
@@
Here $Z_{I}$ is the charge on nucleus $I$, $L_{Y}$ is the number of samples used for fragment $Y$, and $\vec{r}_{k}^{\hskip 1.13809pt(a)}$ is the position of the $k$th unfrozen electron in the $a$th sample. We then evaluate the finite multipole expansion of $\phi_{X}$ (we truncate after the octopole) for use in Eq. (8).
II.4 Overall Cost Scaling
@@
With both the Pauli exclusion errors and electron-electron repulsion errors that arise from freezing electrons now addressed, let us consider the cost of evaluating our energy difference. At each individual sample of the unfrozen electron positions, the local energy $\frac{\hat{H}_{A}\Psi}{\Psi}$ contains $O(1)$ kinetic energy terms. Each costs $O(1)$ time to evaluate if both the orbitals are local and the locality of the Jastrow factor is exploited. Otherwise each costs $O(N)$ time (see Appendix). Summing the $V_{AX}$ interactions requires $O(1)$ time for the standard two-particle Coulomb interactions (note that, for a given spotlight position, $w_{A}$ and $w_{B}$ are only nonzero for $O(1)$ nuclei) and $O(N)$ time for the charge-multipole interactions, although this could be reduced via the fast multipole method in sufficiently large systems. Rokhlin (1985); Greengard and Rokhlin (1987) Thus, if as discussed above the uncertainty contributions from the ordered fragments decay rapidly enough that the total number of samples required across all spotlight positions is $O(1)$, then the overall cost of evaluating the energy difference for our local chemical change will add up to $O(N)$ when nonlocal orbitals are used, and it could become sublinear in large systems if local orbitals, Jastrow locality, and the fast multipole method are employed. However, the cost of first evaluating the multipoles would still be $O(N)$, as we can expect to need $O(1)$ samples for each of the $O(N)$ multipoles.
II.5 Numerical Stability
@@
In this spotlight approach, as in regular VMC, care is required to mitigate the buildup of finite precision arithmetic errors when repeatedly executing low-rank matrix inverse updates via Sherman-Morrison. In fact, there are three contributing factors that make this issue particularly important in the present approach. First, the inverse matrix $(YZ)^{-1}$ on which we would like to perform rank one updates after each one-electron move already contains a matrix inverse inside of it: $Z$ is built from columns of $(M_{\mathrm{init}})^{-1}$. Although we do not update $Z$ as the unfrozen electrons move around, we do update $Y$, and in numerical testing we find that the condition number of $YZ$ can become more extreme than that of the standard $M$ that would be used in regular VMC, leading to faster round-off buildup. Second, in this study at least, we are employing explicitly cusped core orbitals Quady, Bumann, and Neuscamman (2025) whose values can range by multiple orders of magnitude depending on how close to the nucleus an electron is. All else being equal, a larger range of element magnitudes tends to worsen matrix condition numbers (consider, for example, what happens in the case of a diagonal matrix). Third, the space warp transformation we employ in correlated sampling can place the warped electrons near a node even when, pre-warp, the guiding Markov chain is not close to a node. The larger the number of electrons, the more nodes there are to contend with, and we have observed that poor conditioning of the post-warp $YZ$ matrix is more common in larger systems.
@@
We address the first two issues by limiting the number of steps we allow the Markov chain to take before explicitly re-inverting $YZ$ and the third issue by imposing an approximate modification of our probability distribution. Note that, we could, if we wanted, explicitly re-invert $YZ$ at every step while still maintaining the same cost scaling, as $YZ$’s matrix dimension is $O(1)$. However, especially when employing buffer regions, the number of unfrozen electrons is large enough to make a Sherman-Morrison approach worthwhile. The question is where to strike the balance between practical efficiency and numerical stability, and, at least in the systems tested here, we find that re-inverting $YZ$ after each unfrozen electron has attempted 100100100 moves to be an effective compromise. To address our third issue, we note that the conditioning of the pre-warp and post-warp matrices is likely to differ the most when the numerator of the correlated sampling weight factor $W_{i}$ (see Appendix) is significantly different than one. This numerator is connected to the relative condition number, because it contains the ratio of the post-warp to pre-warp determinants inside it. To avoid samples that are likely to be ill conditioned, we simply set the probability distribution equal to zero when this numerator is less than $10^{-5}$ or greater than 101010. This is of course an approximation, but we note that it leads to increasing the Metropolis rejection rate by just 0.01%, and, as we will see in our results, it does not prevent the method from predicting correct energy differences. It does, however, make for a significant improvement in the round-off error of our rank one updates, as desired.
II.6 Initialization

### v2

@@ Abstract
We investigate an approximate sampling scheme that can significantly reduce the cost scaling of variational Monte Carlo when it is employed to predict the energy differences associated with local chemical changes. Inspired by side-chaining and embedding methods, this spotlight sampling approach adopts an approximate fragmented Hamiltonian and correlated sampling to reduce cost scaling to the point that it is essentially linear with system size, with the potential to go sub-linear if certain conditions are met. In tests on bond stretching energies in alcohols, hydrogen dimer chains, and molecules with various degrees of $\pi$-system delocalization, we observe the anticipated linear scaling as well as an explicit cost crossover with standard variational Monte Carlo.
@@
{tocentry}[Uncaptioned image]
1  Introduction
Simultaneously capturing the effects of both weak and strong electron correlation remains challenging for traditional quantum chemistry methods, especially in excited states and large systems. 5, 44, 20, 2, 53, 21, 54, 35 In many difficult cases, Quantum Monte Carlo (QMC) methods have proven to be effective alternatives. 43, 34, 24, 10, 3, 38, 14, 11, 12, 39, 37, 36 For example, in some excited states, even variational Monte Carlo (VMC), perhaps the most straightforward QMC technique, has proven more accurate than full-triples equation of motion coupled cluster theory. 37, 36 However, the significant computational cost required to evaluate large numbers of random samples continues to limit the applicability of both VMC and QMC methods more generally. Indeed, such sampling places an unusually large prefactor in front of VMC’s quartic cost growth with system size, 6 meaning that the advantages of this modest polynomial scaling can be difficult to realize in practice. In this study, we show that a combination of approximate fragmentation and correlated sampling can reduce both cost and scaling while maintaining accuracy when using VMC to predict energy differences associated with local chemical changes, as for example when stretching a chemical bond.
It has long been recognized that, thanks to correlated sampling techniques, it can be advantageous to directly estimate energy differences in VMC rather than attempting to make independent estimates of the energy before and after a chemical change. 51 The basic idea is that, for regions of the system that are only weakly affected by the change, statistical fluctuations will mostly cancel out so long as essentially the sampling of electron positions in those regions is used on either side of the energy difference. The space warp transformation, which we employ in the present study, achieves this goal explicitly. 45, 16 What’s more, by trading the task of estimating two extensive energies for the task of estimating an intensive energy difference, such correlated sampling approaches should in general drop the cost scaling of VMC from quartic to cubic. Here, we seek to combine this advantage with the advantages of local fragmentation approaches, in particular approaches inspired by recent side-chaining methods. 6, 15
Locality has, more generally, long been exploited in attempts to mitigate the cost of VMC. 26, 31, 7, 42 Early works focused on leveraging the sparsity of the Slater matrix when using local orbitals in order to reduce the costs associated with Slater determinant ratios, 56, 30, 1, 4 which can also be mitigated by complimentary delayed update methods. 32 More recently, Feldt and coworkers have shown that that evaluating local side chains, in which only a few of the electrons near a given nucleus are allowed to move, is an effective way to reduce the cost of dealing with core electrons. 15 This side-chaining approach was then generalized in a study by Bienvenu and coworkers, in which they show that it can reduce the cost of VMC sampling by one power of the system size. 6
These side-chaining methods have strong parallels to embedding methods, 25, 47, 55, 57 in that the side chains sample some electrons well while not sampling the “environment” electrons at all, but they differ in one crucial respect. Unlike embedding methods, in which the interaction energy between the system and its environment is approximated, these two side-chain studies carefully set up their overall energy estimates to be unbiased, and so they do not add any approximation. This juxtaposition begs the question: what might we be able to gain if we instead intentionally introduced simplifications that biased the results? Can the cost growth with system size be reduced further while maintaining reasonable accuracy for relevant chemical quantities like bond stretching energies? In this study, we embark on a preliminary investigation of these questions, concluding that, yes, for some types of predictions, a combination of correlated sampling, side-chaining-inspired fragmentation, and embedding-inspired Hamiltonian approximations allow for VMC cost scaling to be reduced further, in principle to the point that it can become linear or even sub-linear in the system size.
Unlike many embedding methods that nestle a high-level method inside a low-level method, such as diffusion Monte Carlo within density functional theory 13 or coupled cluster within density functional theory, 28 the present approach is to use the same level of theory everywhere but to approximate its sampling and long-range electrostatics in order to improve efficiency. Thus, one should not expect accuracy to improve relative to the VMC approach that is being modified. At the same time, however, one need not rely on a lower-level method for any part of the system. At this early stage, we make no attempt to show in which contexts this approach may be preferrable over more traditional embedding approaches. Instead, this study focuses on establishing the basic feasibility of the approach both in principle and in practice and on delineating the conditions under which it should be possible to reach linear and sub-linear scaling.
This paper is organized as follows. We begin in Section 2.1 with a theoretical analysis of how the cost scaling of drawing samples in VMC can be reduced by a combination of various techniques, within which we include a review of the fast Slater determinant math used in Markov chains in which only a small number of electrons move. During this analysis, we do not address the question of accuracy, relying instead on the assumption that certain sampling conditions can be met without doing too much damage to VMC’s energy difference predictions. After completing that analysis, we then introduce in Section 2.2 an explicit approximation scheme, spotlight sampling, which seeks to meet these conditions while maintaining accuracy, focusing in particular on how to deal with both erroneous Pauli exclusion effects, and, in Section 2.3, medium and long range Coulomb interactions. With an approximate Hamiltonian defined, Section 2.4 then completes the cost analysis by showing that its local energy evaluations do not alter the conclusions of the initial sampling cost analysis. We then touch on (Section 2.5) how this approach intersects with the more general VMC challenge of maintaining numerical stability during long chains of low-rank Slater matrix inverse updates before turning to our results. In these results (Section 3), we first verify the necessity of spotlight sampling’s buffering approach, after which we explore proof of principle calculations that demonstrate the cost advantages and test accuracy in a series of alcohols’ O-H bond stretches and in C=C bond stretches in examples with differing degrees of $\pi$ system delocalization. We then end in Section 4 with some concluding remarks.
2  Theory
2.1  Reducing the Cost of Sampling
$\begin{split}E&=\frac{\langle\Psi|\hat{H}|\Psi\rangle}{\langle\Psi|\Psi\rangle}\\ &=\frac{\int\Psi(R)\hat{H}\Psi(R)dR}{\int\Psi(R)\Psi(R)dR}\\ &=\int\rho(R)\frac{\hat{H}\Psi}{\Psi}dR\\ &=\int\rho(R)E_{L}(R)dR\approx\frac{1}{K}\sum_{m}^{K}E_{L}(R_{m})\end{split}$ (1)
Here $E_{L}(R)=\frac{\hat{H}\Psi(R)}{\Psi(R)}$ is the local energy at the electron positions $R_{m}$ of the $m$th sample, and $\rho(R)=|\Psi(R)|^{2}/\int|\Psi(R)|^{2}dR$ is the 3N-dimensional probability distribution for the $N$ electrons. In practice, $\Psi$ is often chosen as $\Psi(R)=e^{J(R)}D(R)$, where $D(R)$ is a Slater determinant and $e^{J(R)}$ is a Jastrow factor,23, 18 while the $K$ samples are drawn from $\rho(R)$ via a Metropolis-Hastings Markov chain 33, 22 in which one electron is moved at a time. 50
@@ 2.1  Reducing the Cost of Sampling
To reduce the cost of working with $M^{-1}$, it is updated at $O(N^{2})$ cost via the Sherman-Morrison formula after each accepted move. 9, 52, 27 Generating a new and independent sample for use in Eq. (1) requires $O(N)$ of these one-electron moves, and, at least when we are attempting to estimate the total energy $E$, the overall number of samples $K$ needs to grow linearly with $N$ to counteract the linear growth in the variance. 6 All told, the computational effort comes out as $O(N^{4})$, which breaks down as taking $O(N)$ samples, each of which requires $O(N)$ moves, each of which bears an $O(N^{2})$ cost for the Sherman-Morrison update.
In this study, we will explore a combination of correlated sampling and fragmentation in order to approximate energy differences for local chemical changes at reduced cost. To start, let us focus first on the effort required to draw samples, and in particular how a fragmentation of the system can reduce this effort. Imagine if, as in a PMC side chain, 6 we only ever had to consider situations in which a small group of O(1) electrons had moved away from their initial positions, while all other electrons remained frozen where they started. In that case, only O(1) of the rows of $M$ would ever change, allowing for an efficient evaluation of a key determinant ratio.
@@
In this study, our goal is more specific than total energies: by also leveraging correlated sampling, we wish to predict energy changes due to local chemical changes at an even lower cost. To do so, we make use of Filippi and Umrigar’s space warp correlated sampling technique. 16 In this approach, whose mathematical details are in the Appendix, electron positions are sampled at one molecular geometry and “warped” to the other geometry in a way that leaves unchanged the positions of electrons near atoms that did not move. We would therefore anticipate that the energy difference $\Delta E^{(l)}$ and energy difference uncertainty $\sigma^{(l)}$ for fragment $l$ will be relatively small if the $l$th fragment’s atoms do not move during the local change, and relatively large if its atoms do move. If we were to sort the fragments in decreasing order of $\sigma^{(l)}$ and could bound these uncertainties by $a/l$ when using some fixed and $O(1)$ fragment sample size $K_{f}$, where $a$ is some constant, then estimating the overall energy difference for the local change would no longer require us to take $O(N)$ samples per fragment. Indeed, $O(1)$ samples per fragment would now suffice. For example, with our bound, $K_{f}$ samples per fragment would give an overall uncertainty of
@@ 2.2  Spotlight Sampling
2.3  Long Range Electrostatics
@@
Here $Z_{I}$ is the charge on nucleus $I$, $L_{Y}$ is the number of samples used for fragment $Y$, and $\vec{r}_{k}^{\hskip 1.13809pt(a)}$ is the position of the $k$th unfrozen electron in the $a$th sample. We then evaluate the finite multipole expansion of $\phi_{X}$ (we truncate after the octupole) for use in Eq. (8).
2.4  Overall Cost Scaling
@@
With both the Pauli exclusion errors and electron-electron repulsion errors that arise from freezing electrons now addressed, let us consider the cost of evaluating our energy difference. At each individual sample of the unfrozen electron positions, the local energy $\frac{\hat{H}_{A}\Psi}{\Psi}$ contains $O(1)$ kinetic energy terms. Each costs $O(1)$ time to evaluate if both the orbitals are local and the locality of the Jastrow factor is exploited. Otherwise each costs $O(N)$ time (see Appendix). Summing the $V_{AX}$ interactions requires $O(1)$ time for the standard two-particle Coulomb interactions (note that, for a given spotlight position, $w_{A}$ and $w_{B}$ are only nonzero for $O(1)$ nuclei) and $O(N)$ time for the charge-multipole interactions, although this could be reduced via the fast multipole method in sufficiently large systems. 41, 19 Thus, if as discussed above the uncertainty contributions from the ordered fragments decay rapidly enough that the total number of samples required across all spotlight positions is $O(1)$, then the overall cost of evaluating the energy difference for our local chemical change will add up to $O(N)$ when nonlocal orbitals are used, and it could become sublinear in large systems if local orbitals, Jastrow locality, and the fast multipole method are employed. However, the cost of first evaluating the multipoles would still be $O(N)$, as we can expect to need $O(1)$ samples for each of the $O(N)$ multipoles.
2.5  Numerical Stability
@@
In this spotlight approach, as in regular VMC, care is required to mitigate the buildup of finite precision arithmetic errors when repeatedly executing low-rank matrix inverse updates via Sherman-Morrison. In fact, there are three contributing factors that make this issue particularly important in the present approach. First, the inverse matrix $(YZ)^{-1}$ on which we would like to perform rank one updates after each one-electron move already contains a matrix inverse inside of it: $Z$ is built from columns of $(M_{\mathrm{init}})^{-1}$. Although we do not update $Z$ as the unfrozen electrons move around, we do update $Y$, and in numerical testing we find that the condition number of $YZ$ can become more extreme than that of the standard $M$ that would be used in regular VMC, leading to faster round-off buildup. Second, in this study at least, we are employing explicitly cusped core orbitals 40 whose values can range by multiple orders of magnitude depending on how close to the nucleus an electron is. All else being equal, a larger range of element magnitudes tends to worsen matrix condition numbers (consider, for example, what happens in the case of a diagonal matrix). Third, the space warp transformation we employ in correlated sampling can place the warped electrons near a node even when, pre-warp, the guiding Markov chain is not close to a node. The larger the number of electrons, the more nodes there are to contend with, and we have observed that poor conditioning of the post-warp $YZ$ matrix is more common in larger systems.
@@
We address the first two issues by limiting the number of steps we allow the Markov chain to take before explicitly re-inverting $YZ$ and the third issue by imposing an approximate modification of our probability distribution. Note that, we could, if we wanted, explicitly re-invert $YZ$ at every step while still maintaining the same cost scaling, as $YZ$’s matrix dimension is $O(1)$. However, especially when employing buffer regions, the number of unfrozen electrons is large enough to make a Sherman-Morrison approach worthwhile. The question is where to strike the balance between practical efficiency and numerical stability, and, at least in the systems tested here, we find that re-inverting $YZ$ after each unfrozen electron has attempted $100$ moves to be an effective compromise. To address our third issue, we note that the conditioning of the pre-warp and post-warp matrices is likely to differ the most when the numerator of the correlated sampling weight factor $W_{i}$ (see Appendix) is significantly different than one. This numerator is connected to the relative condition number, because it contains the ratio of the post-warp to pre-warp determinants inside it. To avoid samples that are likely to be ill conditioned, we simply set the probability distribution equal to zero when this numerator is less than $10^{-5}$ or greater than $10$. This is of course an approximation, but we note that it leads to increasing the Metropolis rejection rate by just 0.01%, and, as we will see in our results, it does not prevent the method from predicting correct energy differences. It does, however, make for a significant improvement in the round-off error of our rank one updates, as desired.
In the future, it will be important to test two ideas related to numerical stability. First, as a system gets larger and larger, more and more of the electrons are far away from the local chemical change and thus are unaffected by the space warp transformation. It may be that the amount of additional difficulty created by the corresponding correlated sampling ratios saturates in large systems, as the positions of the far away electrons should have less and less effect on the probability distribution near the chemical change (this should be true for both the pre-warp and post-warp distribution). If the space warp effect does not saturate with increasing system size, then an alternative correlated sampling approach may be required in large systems. Second, as in most areas of VMC, we expect that using pseudopotentials will improve numerical conditioning. In particular, they will eliminate the large orbital-value fluctuations that come from cusped core orbitals, which in turn should improve the conditioning of both the Slater matrix and our $YZ$ matrix product. On this front, the question would be how big the effect is. Does hiding the core electrons improve numerical stability of spotlight sampling only a little, or does it offer large improvements? Although our current pilot code is unable to answer these questions at present, we flag them as important considerations for future work.
2.6  Initialization